can any9one help me solve this problem. def and abc are complementary angles and abc is three times as large as def. determine the measure of each angles. students must show all work step by stepn to receive full credit.
Answer by ankor@dixie-net.com(22740)

def and abc are complementary angles and abc is three times as large as def.
determine the measure of each angles.
Let x = angle abc
then it's complement, def = (90-x)
:
it says,"abc is three times as large as def. ", therefore
x = 3(90-x)
x = 270 - 3x
x + 3x = 270
now you should be able to solve this